Statement: “If n is any even integer, then ”
Using Direct proof method
1
Expert's answer
2020-11-08T18:36:07-0500

Prove the statement: if n is any even integer, then (−1)n = 1.

Proof:

Suppose n is even integer. Then by the definition of even numbers, n = 2k for some integer k.

Hence, by the laws of exponents of algebra, we have


"(-1)^n=(-1)^{2k}=((-1)^2)^k=(1)^k=1"

This is what was to be shown.

And this completes the proof.